# Adaptive Depth Networks with Skippable Sub-Paths

### v2

@@ Abstract
Predictable adaptation of network depths can be an effective way to control inference latency and meet the resource condition of various devices. However, previous adaptive depth networks do not provide general principles and a formal explanation on why and which layers can be skipped, and, hence, their approaches are hard to be generalized and require long and complex training steps. In this paper, we present a practical approach to adaptive depth networks that is applicable to various networks with minimal training effort. In our approach, every hierarchical residual stage is divided into two sub-paths, and they are trained to acquire different properties through a simple self-distillation strategy. While the first sub-path is essential for hierarchical feature learning, the second one is trained to refine the learned features and minimize performance degradation if it is skipped. Unlike prior adaptive networks, our approach does not train every target sub-network in an iterative manner. At test time, however, we can connect these sub-paths in a combinatorial manner to select sub-networks of various accuracy-efficiency trade-offs from a single network. We provide a formal rationale for why the proposed training method can reduce overall prediction errors while minimizing the impact of skipping sub-paths. We demonstrate the generality and effectiveness of our approach with convolutional neural networks and transformers.
@@ 1 Introduction
Modern deep neural networks such as convolutional neural networks (CNNs) and transformers [1] provide state-of-the-art performance at high computational costs, and, hence, lots of efforts have been made to leverage those inference capabilities in various resource-constrained devices. Those efforts include compact architectures [2, 3], network pruning [4, 5], weight/activation quantization [6], knowledge distillation [7], to name a few. However, those approaches provide static accuracy-efficiency trade-offs, and, hence, it is infeasible to deploy one single model to meet devices with all kinds of resource-constraints.
@@
Figure 1: (a) During training, each residual stage of a network is divided into two sub-paths. The layers in the second sub-paths are optimized to minimize performance degradation even if they are skipped. (b) At test time, we can instantly select $2^{N_{r}}$ parameter-sharing sub-networks from a single network by systematically skipping the last half layers of $N_{r}$ residual stages. (c) The sub-networks selected from a single network form a better Pareto frontier compared to individual networks.
In this work, we introduce an architectural pattern and training method for adaptive depth networks that is generally applicable to various networks, e.g., CNNs and transformers. In the proposed adaptive depth networks, every residual stage is divided into two sub-paths and the sub-paths are trained to have different properties. While the first sub-paths are mandatory for hierarchical feature learning, the second sub-paths are optimized to incur minimal performance degradation even if they are skipped. In order to enforce this property of the second sub-paths, we propose a simple self-distillation strategy, in which only the largest sub-network (or, super-net) and the smallest sub-network (or, base-net) are exclusive used as a teacher and a student, respectively, as shown in Figure-1-(a). The proposed self-distillation strategy does not require exhaustive training of every target sub-network, resulting in significantly shorter training time than prior adaptive networks. Nevertheless, at test time, sub-networks with various depths can be selected instantly from a single network by connecting these sub-paths in a combinatorial manner, as shown in Figure 1-(b). Further, these sub-networks with varying depths outperform individually trained non-adaptive networks due to the regularization effect, as shown in Figure 1-(c).
@@
To authors’ best knowledge, this is the first approach to adaptive networks that provides a general principle and theoretic basis for how depth adaptation can be supported with minimal performance degradation and training effort. This allows us to avoid typical exhaustive training of target sub-networks and instead construct sub-networks of varying depths from specifically trained sub-paths. We anticipate these advances will facilitate future research and applications of adaptive networks.
@@ 2 Related Work
Dynamic Networks: Dynamic networks [21] are another class of adaptive networks that exploit additional control networks or decision gates for input-dependent adaptation of CNN models [22, 23, 24, 25, 26] and transformers [27, 28, 29, 30]. In particular, most dynamic networks for depth-adaptation have some kinds of decision gates at every layers (or blocks) that determine if the layers can be skipped [31, 22, 32, 33]. These approaches are based on the thought that some layers can be skipped on ‘easy’ inputs. However, the learned policy for skipping layers is opaque to users and does not provide a formal description of which layers can be skipped for a given input. Therefore, the network depths cannot be adapted in a predictable manner to meet the resource condition of target devices.
@@ 3 Adaptive Depth Networks
Figure 2: Illustration of a residual stage with two sub-paths. While the first (blue) sub-path is mandatory for hierarchical feature learning, the second (orange) sub-path can be skipped for efficiency. The layers in the skippable sub-path are trained to preserve the feature distribution from $\mathbf{h}^{s}_{base}$ to $\mathbf{h}^{s}_{super}$ using the proposed self-distillation strategy. Having similar feature distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as an input $\mathbf{h}^{s}$ to the next residual stage. In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited when the second sub-path is skipped. This architectural pattern is applicable to both CNNs and transformers.
@@ 3.1 Architectural Pattern for Depth Adaptation
In hierarchical networks, such as ResNets [34] and Swin transformers [38], there are typically 4 to 7 consecutive residual stages. In a network with $N_{r}$ residual stages, the $s$-th ($s=1,...,N_{r}$) residual stage consists of $L$ identical residual blocks  [^1] In transformers, residual blocks are called encoder-decoder blocks. In most networks, $L$ is usually 3 to 6. that transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
@@
To this end, we propose an architectural pattern for adaptive depth networks, in which every residual stage is divided into two sub-paths, or $\mathbf{F}^{s}_{base}$ and $\mathbf{F}^{s}_{skippable}$ as in Equation 1 and Figure 2. We train these two sub-paths to have different properties. While $\mathbf{F}^{s}_{base}$ is trained to learn feature representation $\mathbf{h}^{s}_{base}$ (= $\mathbf{h}^{s}_{L/2+1}$) with no constraint, the second sub-path $\mathbf{F}^{s}_{skippable}$ is constrained to preserve the feature level of $\mathbf{h}^{s}_{base}$ and only refine it to produce $\mathbf{h}^{s}_{super}(=\mathbf{h}^{s}_{L+1}$). Since layers in $\mathbf{F}^{s}_{base}$ perform essential transformations for hierarchical feature learning, they cannot be bypassed during inference. In contrast, layers in $\mathbf{F}^{s}_{skippable}$ can be skipped to save computation since they only refine $\mathbf{h}^{s}_{base}$. If $\mathbf{F}^{s}_{skippable}$ is skipped, then intermediate features $\mathbf{h}^{s}_{base}$ becomes the input to the next residual stage. Therefore, $2^{N_{r}}$ sub-networks with different accuracy-efficiency trade-offs can be selected from a single network by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{r})$ or not.
Table 4 in Appendix A.1 provides details on how this pattern is applied to both CNNs and transformers.
@@ 3.2 Training Sub-Paths with Self-Distillation
In Algorithm 1, the largest and the smallest sub-networks of $\mathbf{M}$, which are called super-net and the base-net, respectively, are exploited. During the forward passes of the super-net and the base-net in steps 5 and 8, respectively, intermediate features $\mathbf{h}_{super}$ and $\mathbf{h}_{base}$ are obtained. In steps 9-10, $D_{KL}(\mathbf{h}^{s}_{super}||\mathbf{h}^{s}_{base})$ in the loss function is explicitly minimized for training data $\mathbf{X}$. Further, this self-distillation step gives the effect of transferring the knowledge from $\mathbf{h}^{s}_{super}$ to $\mathbf{h}^{s}_{base}$ at every residual stage. Therefore, $\mathbf{h}^{s}_{base}$ is expected to learn more compact representation from $\mathbf{h}^{s}_{super}$.
Self-distillation itself has been extensively used in prior adaptive networks [11, 10, 13]. However, their primary goal is to train every target sub-network. For example, on each iteration, sub-networks are randomly sampled from a large search space to act as either teachers or students [11], which results in significantly longer training time. In contrast, our method in Algorithm 1 focuses on training sub-paths by exclusively using the super-net as a teacher and the base-net as a student. By focusing on training sub-paths rather than target sub-networks, the training procedure in Algorithm 1 is significantly simplified. At test time, however, $2^{N_{r}}(=\sum_{k=0}^{N_{r}}{C(N_{r},k)})$ different sub-networks can be constructed instantly by connecting these sub-paths in a combinatorial manner. The effect of this self-distillation strategy is investigated in Section 4.4.
@@ 3.3 Analysis of Skippable Sub-Paths
The high-order terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored if $F_{L}(\mathbf{h}_{L}^{s})$ has a small magnitude. In typical residual networks, however, every layer is trained to learn new features with no constraint, and, hence, there is no guarantee that $F_{L}(\mathbf{h}^{s}_{L})$ have small magnitude. In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the proposed self-distillation strategy, and, hence, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation; see Figure 3 for empirical evidence. If we similarly keep expanding the loss function around $\mathbf{h}^{s}_{j}$ ($j=L/2+1,...,L$) while ignoring high order terms, we obtain the following approximation:
@@
Figure 3: $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at residual blocks. In ours, skippable sub-paths (orange areas) minimally change the distribution of input h
Empirical Analysis: We can estimate how much the distribution of input $\mathbf{h}$ is transformed by residual function $F$ by measuring $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at each residual block. Figure 3 illustrates $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at every residual block in the baseline ResNet50 and our ResNet50-ADN that is trained according to Algorithm 1  [^2] ImageNet validation dataset is used for this experiment.. Our ResNet50-ADN exhibits greater transformation than ResNet50 in the mandatory sub-paths (blue areas) and less transformation in the skippable sub-paths (orange areas). This result demonstrates that our self-distillation strategy in Algorithm 1 effectively trains the skippable sub-paths to minimally change the input distribution. As a result, the blocks in skippable sub-paths can be skipped with less impact on performance.
@@ 4 Experiments
We use four representative residual networks both from CNNs and vision transformers as base models to apply the proposed architecture pattern: MobileNet V2 [35] is a lightweight CNN model, ResNet [34] is a larger CNN model, and ViT [37] and Swin-T [38] are representative vision transformers. All base models except ViT have hierarchical stages, each with 2 $\sim$ 6 residual blocks. So, according to the proposed architectural pattern, every residual stage is evenly divided into 2 sub-paths for depth adaptation. ViT does not define hierarchical stages and all residual encoder blocks have same spatial dimensions. Therefore, we divide 12 encoder blocks into 4 groups, resembling other residual networks, and select the last encoder block of each group as a skippable sub-path. Details are in Table 4 in Appendix A.1.
We use the suffix ‘-ADN’ to denote our adaptive depth networks. Since our adaptive depth networks have many parameter-sharing sub-networks in a single network, we indicate which sub-network is used for evaluation in parenthesis, e.g., ResNet50-ADN (super-net). For our sub-networks, boolean values are also used to indicate in which residual stages the sub-paths are skipped. For example, ResNet50-ADN (base-net) is equivalent to ResNet50-ADN (TTTT).
@@
MbV2-Base 2.99 0.25 70.7
ViT-b/16-ADN (FFFF) 86.59 17.58 81.3
ViT-b/16-ADN (TTTT) 11.76 79.4
@@
Swin-T-ADN (FFFF) 28.30 4.49 81.8
Swin-T-ADN (TTTT) 2.69 78.7
@@ 4.1 ImageNet Classification
Figure 4-(b) shows Pareto frontiers formed by selected sub-networks of our adaptive depth networks; Table 5 in Appendix A.2 shows the performance of all sub-networks. In Figure 4-(b) and Table 1, several state-of-the-art efficient inference methods and dynamic networks are compared with our base-networks. The result demonstrates that our adaptive depth networks show competitive performance across a range of FLOPs. In Figure 4-(b), it should be noted that individual ResNets trained with knowledge distillation has worse performance than individual ResNets. As reported in previous works, successful knowledge distillation requires a patient and long training [52], and straightforward knowledge distillation using ImageNet does not improve the performance of student models [53, 54]. In contrast, our ResNet50-ADN trained with the proposed self-distillation strategy achieves better performance than counterpart ResNets. This demonstrates that the high performance of adaptive depth networks does not simply come from distillation effect.
@@
Swin-T-ADN (TTTT) 2.7G 40% 78.7%
4.2 Training Time
One major advantage of our adaptive depth networks is that they require significantly shorter training time than prior adaptive networks. Figure 5-(a) shows that training our adaptive depth networks take similar time to train two individual networks combined. This is because our training method in Algorithm 1 exclusively exploits only the super-net and the base-net for self-distillation. In contrast, compared adaptive networks require much longer training time than ours since they have to train every target sub-network explicitly  [^3] Although the prior works we compared for training time may seem outdated, they remain relevant and representative because there has been little progress in reducing the training time of adaptive networks.. For example, on every training iteration, AlphaNet [11] randomly samples sub-networks from its search space for self-distillation. While AlphaNet [11] supports significantly more sub-networks, it should be noted that supporting many sub-networks is not our goal. Instead, our objective is to provide better performance Pareto with several useful sub-networks, as demonstrated in Figure 4-(b). For example, although ResNet50-ADN can select $2^{4}$ sub-networks, Table 5 (in Appendix A.2) shows that some sub-networks are more useful than the others; results in Table 5 show that skipping sub-paths in deeper residual stages is more detrimental to the performance in general, implying that refinement of features is more needed for deeper residual stages.

### v3

@@ Abstract
Predictable adaptation of network depths can be an effective way to control inference latency and meet the resource condition of various devices. However, previous adaptive depth networks do not provide general principles and a formal explanation on why and which layers can be skipped, and, hence, their approaches are hard to be generalized and require long and complex training steps. In this paper, we present a practical approach to adaptive depth networks that is applicable to both convolutional neural networks (CNNs) and transformers with minimal training effort. In our approach, every hierarchical residual stage is divided into two sub-paths, and they are trained to acquire different properties through a simple self-distillation strategy. While the first sub-path is essential for hierarchical feature learning, the second one is trained to refine the learned features and minimize performance degradation even if it is skipped. Unlike prior adaptive networks, our approach does not train every target sub-network exhaustively. At test time, however, we can connect these sub-paths in a combinatorial manner to select sub-networks of various accuracy-efficiency trade-offs from a single network. We provide a formal rationale for why the proposed training method can reduce overall prediction errors while minimizing the impact of skipping sub-paths. We demonstrate the generality and effectiveness of our approach with both CNNs and transformers. Source codes are available at https://github.com/wchkang/depth
@@ 1 Introduction
Modern deep neural networks such as CNNs and transformers [1] provide state-of-the-art performance at high computational costs, and, hence, lots of efforts have been made to leverage those inference capabilities in various resource-constrained devices. Those efforts include compact architectures [2, 3], network pruning [4, 5], weight/activation quantization [6], knowledge distillation [7], to name a few. However, those approaches provide static accuracy-efficiency trade-offs, and, hence, it is infeasible to deploy one single model to meet devices with all kinds of resource-constraints.
@@
Figure 1: (a) During training, every residual stage of a network is divided into two sub-paths. The layers in every second (orange) sub-path are optimized to minimize performance degradation even if they are skipped. (b) At test time, these second sub-paths can be skipped in a combinatorial manner, allowing instant selection of various parameter sharing sub-networks. (c) The sub-networks selected from a single network form a better Pareto frontier than counterpart individual networks.
In this work, we introduce an architectural pattern and training method for adaptive depth networks that is generally applicable to various networks, e.g., CNNs and transformers. In the proposed adaptive depth networks, every residual stage is divided into two sub-paths and the sub-paths are trained to have different properties. While the first sub-paths are mandatory for hierarchical feature learning, the second sub-paths are optimized to incur minimal performance degradation even if they are skipped. In order to enforce this property of the second sub-paths, we propose a simple self-distillation strategy, in which only the largest sub-network (or, super-net) and the smallest sub-network (or, base-net) are exclusive used as a teacher and a student, respectively, as shown in Figure-1-(a). The proposed self-distillation strategy does not require exhaustive training of every target sub-network, resulting in significantly shorter training time than prior adaptive networks. However, at test time, sub-networks with various depths can be selected instantly from a single network by connecting these sub-paths in a combinatorial manner, as shown in Figure 1-(b). Further, these sub-networks with varying depths outperform individually trained non-adaptive networks due to the regularization effect, as shown in Figure 1-(c).
@@
Our approach uniquely introduces a principle for training selected sub-paths to be skippable with minimal performance loss. This principle allows us to avoid typical exhaustive training of target sub-networks and instead construct sub-networks of varying depths from specifically trained sub-paths. We anticipate these advances will facilitate future research and applications of adaptive networks.
@@ 2 Related Work
Dynamic Networks: Dynamic networks [21] are another class of adaptive networks that exploit additional control networks or decision gates for input-dependent adaptation of CNN models [22, 23, 24, 25, 26] and transformers [27, 28, 29, 30]. In particular, most dynamic networks for depth-adaptation have some kinds of decision gates at every layer (or block) that determine if the layer can be skipped [31, 22, 32, 33]. These approaches are based on the thought that some layers can be skipped on ‘easy’ inputs. However, the learned policy for skipping layers is opaque to users and does not provide a formal description of which layers can be skipped for a given input. Therefore, the network depth cannot be controlled by users in a predictable manner to meet the resource condition of target devices.
@@ 3 Adaptive Depth Networks
Figure 2: Illustration of a residual stage with two sub-paths. While the first (blue) sub-path is mandatory for hierarchical feature learning, the second (orange) sub-path can be skipped for efficiency. The layers in the skippable sub-path are trained to preserve the feature distribution from $\mathbf{h}^{s}_{base}$ to $\mathbf{h}^{s}_{super}$ using the proposed self-distillation strategy. Having similar distributions, either $\mathbf{h}^{s}_{base}$ or $\mathbf{h}^{s}_{super}$ can be provided as input $\mathbf{h}^{s}$ to the next residual stage. In the mandatory sub-path, another set of batch normalization operators, called skip-aware BNs, is exploited if the second sub-path is skipped. These sub-paths are building blocks to construct sub-networks of varying depths.
@@ 3.1 Architectural Pattern for Depth Adaptation
In modern hierarchical networks, such as ResNets [34] and Swin transformers [38], there are typically 4 to 7 consecutive residual stages. [^1] For vision transformers, residual blocks/stages refer to encoder blocks/stages. In a network with $N_{r}$ stages, the $s$-th ($s=1,...,N_{r}$) stage consists of $L$ (typically $L\in[3,6]$) identical residual blocks that transform input features $\mathbf{h}^{s}_{1}$ additively to produce the output features $\mathbf{h}^{s}$, as follows:
@@
To this end, we propose an architectural pattern for adaptive depth networks, in which every residual stage is divided into two sub-paths, or $\mathbf{F}^{s}_{base}$ and $\mathbf{F}^{s}_{skippable}$ as in Equation 1 and Figure 2. We train these two sub-paths to have different properties (Section 3.2). While $\mathbf{F}^{s}_{base}$ is trained to learn feature representation $\mathbf{h}^{s}_{base}$ (= $\mathbf{h}^{s}_{L/2+1}$) with no constraint, the second sub-path $\mathbf{F}^{s}_{skippable}$ is constrained to preserve the feature level of $\mathbf{h}^{s}_{base}$ and only refine it to produce $\mathbf{h}^{s}_{super}(=\mathbf{h}^{s}_{L+1}$). Since layers in $\mathbf{F}^{s}_{base}$ perform essential transformations for hierarchical feature learning, they cannot be bypassed during inference. In contrast, layers in $\mathbf{F}^{s}_{skippable}$ can be skipped for efficiency since they only refine $\mathbf{h}^{s}_{base}$. If $\mathbf{F}^{s}_{skippable}$ is skipped, then $\mathbf{h}^{s}_{base}$ becomes the input to the next residual stage. Therefore, $2^{N_{r}}$ (=$\sum_{k=0}^{N_{r}}{C(N_{r},k)}$) sub-networks with varying accuracy-efficiency trade-offs can be selected from a single network by choosing whether to skip $\mathbf{F}^{s}_{skippable}(s=1,...,N_{r})$ or not (Table 5).
This architectural pattern is agnostic to the type of residual blocks. The residual function $F$ can be convolution layers for CNNs and self-attention + MLP layers for transformers. Table 4 in Appendix A.1 provides details on how this pattern is applied to both CNNs and transformers.
@@ 3.2 Training Sub-Paths with Self-Distillation
In Algorithm 1, the largest and the smallest sub-networks of $\mathbf{M}$, which are called super-net and the base-net, respectively, are exploited. In steps 9-10, the hierarchical features (and outputs) from the super-net and the base-net are encouraged to have similar distributions through self-distillation. Due to the architectural pattern of interleaving the mandatory and the skippable sub-paths, minimizing $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ also minimizes $D_{KL}(\mathbf{h}_{super}\|\mathbf{h}_{base})$ implicitly. Therefore, only $D_{KL}(\mathbf{\hat{y}}_{super}\|\mathbf{\hat{y}}_{base})$ can be included in $loss_{base}$ if the extraction of intermediate features is tricky.
Self-distillation has been extensively used in prior adaptive networks [11, 10, 13]. However, their primary goal is to train every target sub-network. For example, on each iteration, sub-networks are randomly sampled from a large search space to act as either teachers or students [11], which takes significantly longer training time and can generate conflicting training objectives for different parameter-sharing sub-networks [15, 16]. In contrast, our method in Algorithm 1 focuses on training sub-paths by exclusively using the super-net as a teacher and the base-net as a student. By focusing on sub-paths, the training procedure in Algorithm 1 is significantly simplified and avoids potential conflicts among sub-networks. At test time, however, various sub-networks can be selected by connecting these sub-paths. The effect of this self-distillation strategy is investigated in Section 4.4.
@@ 3.3 Analysis of Skippable Sub-Paths
The high-order terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored if $F_{L}(\mathbf{h}_{L}^{s})$ has a small magnitude. In typical residual networks, however, every layer is trained to learn new features with no constraint, and, hence, there is no guarantee that $F_{L}(\mathbf{h}^{s}_{L})$ have small magnitude. In contrast, in our adaptive depth networks, the residuals in $\mathbf{F}^{s}_{skippable}$ are explicitly enforced to have small magnitude through the proposed self-distillation strategy (refer to Figure 3 for empirical evidence). As a result, the terms in $\mathcal{O}(F_{L}(\mathbf{h}_{L}^{s}))$ can be ignored for the approximation. If we similarly keep expanding the loss function around $\mathbf{h}^{s}_{j}$ ($j=L/2+1,...,L$) while ignoring high order terms, we obtain the following approximation:
@@
Figure 3: $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at residual blocks. In ours, skippable sub-paths (orange areas) minimally change the distribution of input h.
Empirical Analysis: We can estimate how much the distribution of input $\mathbf{h}$ is transformed by residual function $F$ by measuring $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at each residual block. Figure 3 illustrates $||F(\mathbf{h})||_{2}/||\mathbf{h}||_{2}$ at every residual block in the baseline ResNet50 and our ResNet50-ADN that is trained according to Algorithm 1.222ImageNet validation dataset is used for this experiment. Our ResNet50-ADN exhibits greater transformation than ResNet50 in the mandatory sub-paths (blue areas) and less transformation in the skippable sub-paths (orange areas). This result demonstrates that our self-distillation strategy in Algorithm 1 effectively trains the skippable sub-paths to minimally change the input distribution. As a result, the blocks in skippable sub-paths can be skipped with less impact on performance.
@@ 4 Experiments
We use networks both from CNNs and vision transformers as base models to apply the proposed architecture pattern: MobileNet V2 [35] is a lightweight CNN model, ResNet [34] is a larger CNN model, and ViT [37] and Swin-T [38] are representative vision transformers. All base models except ViT have hierarchical stages, each with 2 $\sim$ 6 residual blocks. So, according to the proposed architectural pattern, every stage is evenly divided into 2 sub-paths for depth adaptation. ViT does not define hierarchical stages and all residual encoder blocks have same spatial dimensions. Therefore, we divide 12 encoder blocks into 4 groups, resembling other residual networks, and select the last encoder block of each group as a skippable sub-path. Details are in Table 4 in Appendix A.1. These models are trained according to Algorithm 1. For self-distillation, only the final outputs from the super-net and the base-net, or $\mathbf{\hat{y}}_{super}$ and $\mathbf{\hat{y}}_{base}$ respectively, are used since exploiting intermediate features for explicit self-distillation has a marginal impact on performance (Section 4.4).
We use the suffix ‘-ADN’ to denote our adaptive depth networks. A series of boolean values in parentheses denotes a specific sub-network used for evaluation; each boolean value represents the residual stage where its skippable sub-path is skipped. For example, ResNet50-ADN (FFFF) and ResNet50-ADN (TTTT) correspond to the super-net and the base-net of ResNet50-ADN, respectively.
@@
MbV2-Base 2.98 0.22 70.2
ViT-b/16-ADN (FFFF) 86.59 17.58 81.4
ViT-b/16-ADN (TTTT) 11.76 80.6
@@
Swin-T-ADN (FFFF) 28.30 4.49 81.6
Swin-T-ADN (TTTT) 2.34 78.0
@@ 4.1 ImageNet Classification
Figure 4-(b) shows Pareto frontiers formed by selected sub-networks of our adaptive depth networks; Table 5 in Appendix A.2 shows the performance of all sub-networks. In Figure 4-(b) and Table 1, several state-of-the-art efficient inference methods and dynamic networks are compared with our base-networks. The result demonstrates that our adaptive depth networks show comparable performance across a range of FLOPs. In Figure 4-(b), it should be noted that individual ResNets trained with knowledge distillation has worse performance than individual ResNets. As reported in previous works, successful knowledge distillation requires a patient and long training [52], and straightforward knowledge distillation using ImageNet does not improve the performance of student models [53, 54]. In contrast, our ResNet50-ADN trained with the proposed self-distillation strategy achieves better performance than counterpart ResNets. This demonstrates that the high performance of adaptive depth networks does not simply come from distillation effect.
@@
Swin-T-ADN (TTTT) 2.3G 48% 78.0%
4.2 Training Time and Sub-Networks
One of the key advantages of our training method in Algorithm 1 is its simplicity and significantly reduced training effort compared to previous adaptive networks. Figure 5-(a) shows that the training time of our adaptive depth networks are similar to the training time of two individual networks combined. This is because our training method trains sub-paths, rather than sub-networks, by exploiting only the super-net and the base-net. In contrast, the compared adaptive networks require much longer training time than ours since they have to explicitly apply self-distillation to all target sub-networks. For example, on every training iteration, AlphaNet [11] randomly samples sub-networks from its search space for self-distillation. Although the prior works we compared may seem outdated, they remain relevant and representative because there has been little progress in improving the training cost of adaptive networks.
While AlphaNet [11] supports significantly more sub-networks, it should be noted that supporting many sub-networks is not our goal. Instead, our objective is to provide better performance Pareto with several useful sub-networks, as demonstrated in Figure 4-(b). For example, although ResNet50-ADN has $2^{4}$ sub-networks in its search space, Table 5 in Appendix A.2 demonstrates that some sub-networks manifest better performance than the others even if they have similar depths. The results show that skipping sub-paths in deeper residual stages generally has a more detrimental impact on performance, suggesting that feature refinement is more crucial for deeper residual stages.